2) a. F(s) = S+3 (s+3)2 + 16 b. F(s) = 1 (s-2)2+4 C. F(S) 1 (s-4) d. F(s) = 6-45

Find the Inverse Laplace Transform of the following functions

Please show all work.Thank You.
Answer #1

FSJ = S+3 3 416 [ {FCS-a)} = eat f(t) S43 너는 (St3P 1 16 너 3= 매 2 Cosat 너 is cl } = (-3t cos (47) L{ 83 +34 +16El fear Plays=0ye? + Clear 2.0 IS) (62)2 + 4 {F(S-2)} – cat flt) It (8-22 tu ay 22.3 cm x 13.02.13-2) Ź ezt sinat. 20 F (s)-